# Sensor-Based Distributionally Robust Control for Safe Robot Navigation in Dynamic Environments

### v1

@@ Abstract
We introduce a novel method for safe mobile robot navigation in dynamic, unknown environments, utilizing onboard sensing to impose safety constraints without the need for accurate map reconstruction. Traditional methods typically rely on detailed map information to synthesize safe stabilizing controls for mobile robots, which can be computationally demanding and less effective, particularly in dynamic operational conditions. By leveraging recent advances in distributionally robust optimization, we develop a distributionally robust control barrier function (DR-CBF) constraint that directly processes range sensor data to impose safety constraints. Coupling this with a control Lyapunov function (CLF) for path tracking, we demonstrate that our CLF-DR-CBF control synthesis method achieves safe, efficient, and robust navigation in uncertain dynamic environments. We demonstrate the effectiveness of our approach in simulated and real autonomous robot navigation experiments, marking a substantial advancement in real-time safety guarantees for mobile robots.
@@ 1 Introduction
Ensuring real-time, high-frequency robot control with safety guarantees in dynamic and unstructured environments is crucial for the effective deployment of autonomous mobile robots. Khatib (1986) introduced the seminar artificial potential fields approach to enable collision avoidance for real-time control of mobile robots. This concept has inspired a wealth of research into the joint consideration of path planning and control, including techniques utilizing navigation functions (Rimon and Koditschek, 1992), dynamic windows (Fox et al., 1997), and velocity vector fields (De Lima and Pereira, 2013). A common thread in many of these works is the reliance on accurate map representations (Oleynikova et al., 2017; Herbert et al., 2017; Arslan and Koditschek, 2019; Axelrod et al., 2018; Li et al., 2023b) updated from onboard sensing to facilitate safe autonomous navigation. However, creating accurate maps in real-time with limited on-board computing resources is particularly challenging in dynamic environments. This work addresses this challenge by developing a distributionally robust formulation for safety constraints that directly processes range sensor data, avoiding the need for fast map updates and offering an adaptable solution for safe navigation in unstructured and dynamic environments.
@@
Certificate functions have been introduced as powerful tools for asserting properties of dynamical systems, such as stability, safety, and robustness against uncertainties. Among these, Lyapunov functions (Artstein, 1983; Sontag, 1989) guarantee asymptotic stability for dynamical systems, and barrier functions (Prajna and Jadbabaie, 2004) certify forward invariance for desired safe sets. In recent years, control barrier functions (CBFs) have marked a significant advancement in encoding safety constraints for dynamical systems. In conjunction with control Lyapunov functions (CLFs) for stability guarantees, safe and stable controls can be synthesized online for control-affine nonlinear systems via quadratic programming (QP) (Ames et al., 2019). The CLF-CBF QP framework has become a mainstream approach for synthesizing reliable and safe controls in real-time for various robot systems (Desai and Ghaffari, 2022; Choi et al., 2023; Li et al., 2023a; Liu et al., 2023).
@@
In this paper, we develop a novel approach to impose safety constraints for robot navigation in unknown and dynamic environments. Leveraging distributionally robust optimization (DRO), our method circumvents the requirement of a precisely known occupancy map and deals effectively with the uncertainty in direct range measurements. Conventionally, safe control synthesis with CBFs necessitates perfect knowledge of both the system model and the associated CBF (Ames et al., 2017). Although recent efforts have focused on estimating CBFs using sensory data like LiDAR and RGB-D in unknown environments (Long et al., 2022; Xiao et al., 2022; Abdi et al., 2023; Hamdipoor et al., 2023), these methods involve processing the sensor data to reconstruct the environment, a process that poses significant computational challenges in real-time robot applications.
Our approach addresses this bottleneck by enabling the evaluation of safety constraints at comparable frequency to control. Instead of estimating a CBF constraint and its uncertainty precisely, our formulation enables direct use of the measurements as noisy CBF samples defining a DRO safety constraint. Our formulation enhances the robustness of safe control synthesis in complex real-world scenarios (e.g., thin chair legs, netting in Fig. 1), where constructing accurate CBFs from sensor measurements is particularly challenging. This advancement represents a critical step in ensuring safer and more reliable autonomous robot navigation in dynamically changing and challenging environments.
The main contributions of this paper are as follows.
@@
  • We formulate a novel distributionally robust control barrier function (DR-CBF) constraint directly using onboard range measurements to synthesize safe controls that circumvent the need for fast, precise map updates in dynamic environments.
@@
  • We validate the safety, efficiency, and robustness of our approach in simulated and real experiments of autonomous differential-drive robot navigation in unknown and dynamic environments.
  • We provide an open-source implementation of our proposed CLF-DR-CBF controller [^1] Project page: https://existentialrobotics.org/DR_Safe_Navigation_Webpage/.
@@
Figure 1: ClearPath Jackal robot equipped with a LiDAR sensor navigating in unknown indoor and outdoor environments.
@@ 2 Related Work
Dynamic obstacle avoidance. Robot motion planning algorithms have a rich history, dating back to the 1950s with the introduction of Dijkstra algorithm (Dijkstra, 1959) and $A^{*}$ (Hart et al., 1968) for search-based planning in known environments. Since then, a substantial amount of research has been dedicated to algorithms for collision-free path planning and low-level control that allow robots to follow the planned paths (Lozano-Perez, 1983; Brooks, 1983). A significant contribution was made by Khatib (1986), who introduced artificial potential fields to enable collision avoidance during not only the motion planning stage but also the real-time control of a mobile robot. The formulation was extended to a virtual force field (Borenstein et al., 1991), facilitating safe navigation in uncertain environments. Later, Rimon and Koditschek (1992) developed navigation functions, a particular form of artificial potential functions, which simultaneously ensure collision avoidance and stabilization to a goal configuration. The dynamic window concept was introduced by Fox et al. (1997) to handle dynamic obstacles by proactively filtering out unsafe control actions. Later, this idea was combined with a global desired velocity vector field to allow safe navigation in De Lima and Pereira (2013). More recently, Herbert et al. (2017) developed FaSTrack, a modular framework that utilizes Hamilton-Jacobi reachability analysis and a game-theoretic formulation between planner and controller to enable rapid and safe robot navigation in complex environments. Majd et al. (2021) integrate time-based rapidly-exploring random trees (RRTs) with CBFs to enable safe navigation in dynamic environments densely populated by pedestrians. Recent advancements in reinforcement learning have also led to real-time solutions for obstacle avoidance and navigation (Pfeiffer et al., 2018; Everett et al., 2021; Chen et al., 2022), typically framing the problem as a partially observable Markov decision processe to facilitate model-free learning of control policies.
@@
While these methods show promise, most require knowledge or estimation of environment geometry, typically through topological (Dudek et al., 1978) and metric (Chatila and Laumond, 1985; Borenstein et al., 1991) map representations. The signed distance function (SDF) has emerged as a particularly valuable tool in this regard (Oleynikova et al., 2017; Han et al., 2019; Wu et al., 2023), offering distance and gradient information for safe control and navigation. However, constructing real-time occupancy or SDF estimates is challenging with onboard sensing and limited computing resources, particularly in complex and dynamic environments.
@@
Distributionally robust optimization. Distributionally Robust Optimization (DRO) considers parameter uncertainty in optimization problems, and is particularly effective when dealing with a limited number of uncertainty samples. This method compensates for the potential discrepancy between empirical and true uncertainty distributions, utilizing uncertainty descriptors like moment ambiguity sets (Van Parys et al., 2016), Kullback–Leibler ambiguity sets (Jiang and Guan, 2016), and Wasserstein ambiguity sets (Esfahani and Kuhn, 2018; Xie, 2021). The DRO framework has been increasingly utilized for its robust performance guarantees against distributional uncertainty in control (Boskos et al., 2024; Mestres et al., 2024; Long et al., 2023a; Chriat and Sun, 2023) and robotics (Ren and Majumdar, 2022; Coulson et al., 2021) applications.
@@
Lathrop et al. (2021) proposed the Wasserstein safe RRT, a path planning algorithm offering finite-sample probabilistic safety guarantees in uncertain non-convex obstacle environments. A DRO-based approach was introduced by Ren and Majumdar (2022) to enhance policy robustness by iteratively training with adversarial environments generated through a learned generative model. Long et al. (2023b) proposed a DRO formulation for safe stabilizing control under model uncertainty, assuming that nominal safety and stability certificates are provided. Hakobyan and Yang (2022) introduced a distributionally robust risk map as a safety specification tool for mobile robots, effectively reformulating the optimization problem over an infinite-dimensional probability distribution space into a tractable semidefinite program. Boskos et al. (2023) proposed a distributionally robust coverage control algorithm for a team of robots to optimally deploy in a spatial region with an unknown event probability density. A Wasserstein tube MPC is presented in Aolaritei et al. (2023) for stochastic systems, which utilizes Wasserstein ambiguity sets to construct tubes around nominal trajectories, enhancing robustness and efficiency with limited noise samples available. Ramesh et al. (2023) introduced a model-based approach that utilizes the maximum variance reduction algorithm to efficiently learn nominal transition dynamics and near-optimal distributionally robust policies for stochastic systems.
@@
Safe stabilizing control. Quadratic programming that integrates CLF and CBF constraints presents a certifiable and efficient framework for synthesizing safe stabilizing controls in a variety of tasks. These tasks span from multi-robot safe navigation (Zhang et al., 2023), to safe locomotion of legged robots (Grandia et al., 2021), and to safe humanoid operation (Khazoom et al., 2022). Despite the efficiency of the CLF-CBF QP in synthesizing safe stabilizing controls, it typically relies on perfect knowledge of system dynamics, state estimates, and barrier function constraints. However, in many robotics applications, various sources of uncertainty can significantly impact performance and reliability. Some recent studies have begun to explore this area, to address uncertainty in system dynamics (Dhiman et al., 2023; Emam et al., 2022), state estimates (Daş and Murray, 2022; Wang and Xu, 2023), and barrier function constraints (Long et al., 2021; Hamdipoor et al., 2023). These approaches typically utilize robust and probabilistic models to integrate uncertainty into the QP formulation, leading to convex reformulations that enhance robustness.
@@
Particularly in safe robot navigation, a robot might need to estimate barrier functions based on its (noisy) observations. Long et al. (2021) introduced an incremental online learning approach for estimating the barrier function from LiDAR data and proposed a robust reformulation of the CLF-CBF QP by incorporating estimation errors. Dawson et al. (2022) developed an approach to learn observation-space CBFs utilizing distance data and proposed a two-mode hybrid controller to avoid deadlocks in navigation. A reactive planning algorithm was presented in Liu et al. (2023) for the safe operation of a bipedal robot with multiple obstacles, utilizing a single differentiable CBF derived from LiDAR point clouds. Abdi et al. (2023) proposed a method for learning vision-based CBF from RGB-D images with pre-training, enabling safe navigation of autonomous vehicles in unseen environments. By effectively decomposing and predicting the spatial interactions of multiple obstacles, Yu et al. (2023) proposed compositional learning of sequential CBFs, enabling obstacle avoidance in dense dynamic environments. Keyumarsi et al. (2024) introduced an efficient Gaussian Process-based method for synthesizing CBFs from LiDAR data, showcasing its effectiveness in a turtlebot navigation task. Zhang et al. (2024) introduced an efficient LiDAR-based framework for goal-seeking and exploration of mobile robots in dynamic environments, utilizing minimum bounding ellipses to represent obstacles and Kalman filters to estimate their velocities.
@@
A related line of research to our work is the application of conformal prediction in robot navigation tasks. Conformal prediction (Shafer and Vovk, 2008; Zhao et al., 2024) is a statistical tool for uncertainty quantification that provides valid prediction regions with a user-specified risk tolerance, making it particularly useful for ensuring safety in dynamic environments. Several recent works have explored the integration of conformal prediction into motion planning and control frameworks. Lindemann et al. (2023) used conformal prediction to obtain prediction regions for a model predictive controller. Yang et al. (2023) employed conformal prediction to quantify state estimation uncertainty and design a robust CBF controller based on the estimated uncertainty. An adaptive conformal prediction algorithm was developed by Dixit et al. (2023) to dynamically quantify prediction uncertainty and plan probabilistically safe paths around dynamic agents.
3 Background
@@ 3.2 CLF-CBF Quadratic Program
where $\delta$ denotes a slack variable that relaxes the CLF constraint to ensure feasibility of the QP, controlled by the scaling factor $\lambda>0$. As discussed in Ames et al. (2017); Mestres et al. (2023), if the CBF $h$ has relative degree of $1$ with respect to the system dynamics in (1), the controller in (6) is Lipschitz continuous in $\mathbf{x}$ and piecewise continuous in $t$. This guarantees unique solutions for the closed-loop system, ensuring the safe set ${\cal C}(t)$ remains forward invariant.

### v2

@@ Abstract
We introduce a novel method for mobile robot navigation in dynamic, unknown environments, leveraging onboard sensing and distributionally robust optimization to impose probabilistic safety constraints. Our method introduces a distributionally robust control barrier function (DR-CBF) that directly integrates noisy sensor measurements and state estimates to define safety constraints. This approach is applicable to a wide range of control-affine dynamics, generalizable to robots with complex geometries, and capable of operating at real-time control frequencies. Coupled with a control Lyapunov function (CLF) for path following, the proposed CLF-DR-CBF control synthesis method achieves safe, robust, and efficient navigation in challenging environments. We demonstrate the effectiveness and robustness of our approach for safe autonomous navigation under uncertainty in simulations and real-world experiments with differential-drive robots.
@@ 1 Introduction
Ensuring real-time, high-frequency robot control with safety guarantees in dynamic and unstructured environments is crucial for the effective deployment of autonomous mobile robots. Khatib (1986) introduced the seminal artificial potential fields approach to enable collision avoidance for real-time control of mobile robots. This concept has inspired a wealth of research into the joint consideration of path planning and control, including techniques utilizing navigation functions (Rimon and Koditschek, 1992), dynamic windows (Fox et al., 1997), and velocity vector fields (De Lima and Pereira, 2013). Many work on joint planning and control rely on accurate map representations (Oleynikova et al., 2017; Herbert et al., 2017; Arslan and Koditschek, 2019; Brito et al., 2019; Schaefer et al., 2021; Kondo et al., 2023; Liu et al., 2023a) updated from onboard sensing to facilitate safe autonomous navigation. However, in practice, high-level planning operates at a low frequency and depends on up-to-date maps, while low-level control requires high-frequency updates (Fox et al., 1997; Corke et al., 2000; Frazzoli et al., 2002; Huang and Grizzle, 2023), relying on raw sensor observations due to computational constraints.
@@
Certificate functions have been introduced as powerful tools to assert properties of dynamical systems, such as stability, safety, and robustness against uncertainties. Among these, Lyapunov functions (Artstein, 1983; Sontag, 1989) guarantee asymptotic stability for dynamical systems, while barrier functions (Prajna and Jadbabaie, 2004) certify forward invariance for desired safe sets. In recent years, control barrier functions (CBFs) have marked a significant advancement in encoding safety constraints for dynamical systems. In conjunction with control Lyapunov functions (CLFs) for stability guarantees, safe and stable controllers can be synthesized online for control-affine systems via quadratic programming (QP) (Ames et al., 2019). The CLF-CBF QP framework has become a mainstream approach for synthesizing safe and stable controls in real-time for various robot systems (Desai and Ghaffari, 2022; Choi et al., 2023; Li et al., 2023; Liu et al., 2023b).
@@
In this paper, we develop a novel approach leveraging distributionally robust optimization (DRO) to impose safety constraints for robot navigation in unknown and dynamic environments under measurement and state estimation uncertainty. Conventionally, safe control synthesis with CBFs necessitates perfect knowledge of the robot’s state and model along with the associated CBF (Ames et al., 2017). Although recent efforts have focused on estimating CBFs using onboard measurements in unknown environments (Long et al., 2022; Xiao et al., 2022; Abdi et al., 2023; Hamdipoor et al., 2023; Abuaish et al., 2023), these methods rely on perfect localization and simplified robot shapes. Moreover, they involve processing sensor data to reconstruct the environment, which introduces computational challenges for real-time robot operation.
Our approach allows evaluation of safety constraints at the control frequency. Instead of estimating a CBF constraint and its uncertainty precisely, our formulation uses sensor measurements and state estimates directly as noisy CBF samples to define DRO safety constraints. Our formulation enhances the robustness of safe control synthesis in complex, dynamic real-world scenarios, addressing the challenges posed by state estimation errors and uncertain sensor measurements, which can render theoretically valid CBFs unsafe in practice. The main contributions of this paper are as follows.
@@
  • We formulate a novel distributionally robust control barrier function (DR-CBF) constraint directly using onboard distance measurements to synthesize safe controls for robots with non-trivial shapes.
  • We provide valid probabilistic safety guarantees, ensuring that the robot is safe with a user-defined probability, even under uncertainty in sensor measurements and robot state estimation.
@@
  • We validate the safety, efficiency, and robustness of our approach in simulated and real experiments with autonomous differential-drive robots navigating in unknown and dynamic environments, illustrated in Fig. 1. An open-source implementation of our CLF-DR-CBF controller is available on our project page [^1] Project page: https://existentialrobotics.org/DRO_Safe_Navigation/.
@@
Figure 1: ClearPath Jackal robot equipped with a LiDAR sensor navigating in unknown environments.
@@ 2 Related Work
Dynamic obstacle avoidance. Robot motion planning algorithms have a rich history, dating back to the 1950s with the introduction of the Dijkstra (Dijkstra, 1959) and $A^{*}$ (Hart et al., 1968) algorithms for search-based planning. Since then, a substantial amount of research has been dedicated to algorithms for motion planning and trajectory tracking control (Lozano-Perez, 1983; Corke et al., 2000; Frazzoli et al., 2002; Huang and Grizzle, 2023). A significant contribution was made by Khatib (1986), who introduced artificial potential fields to enable collision avoidance during not only the motion planning stage but also the real-time control of a mobile robot. The formulation was extended to a virtual force field (Borenstein et al., 1991), facilitating safe navigation in uncertain environments. Rimon and Koditschek (1992) developed navigation functions, a particular form of artificial potential functions, which simultaneously ensure collision avoidance and stabilization to a goal configuration. The dynamic window concept was introduced by Fox et al. (1997) to handle dynamic obstacles by proactively filtering out unsafe control actions. More recently, Bajcsy et al. (2019) proposed leveraging Hamilton-Jacobi reachability for autonomous navigation to provide provable safety guarantees in unknown static environments, enabling real-time updates from sensor measurements while accommodating general nonlinear dynamics and planners. Recent advancements in safe motion planning for dynamic environments handle various types of uncertainty, including obstacle locations (Liu et al., 2023a), human behavior prediction (Schaefer et al., 2021) and localization errors (Summers, 2018). Recent advancements in reinforcement learning have also led to real-time techniques for autonomous navigation (Pfeiffer et al., 2018; Everett et al., 2021; Chen et al., 2022), typically framing the problem as a partially observable Markov decision processes to facilitate model-free learning of control policies.
@@
Many of these methods require knowledge or estimation of the environment geometry and dynamics, typically in the form of topological (Dudek et al., 1978) or metric (Chatila and Laumond, 1985; Borenstein et al., 1991) map representations. The signed distance function (SDF) has emerged as a particularly valuable tool in this regard (Oleynikova et al., 2017; Han et al., 2019; Wu et al., 2023), offering distance and gradient information for safe control and navigation. However, constructing occupancy or SDF maps at control frequency is challenging in complex and dynamic environments.
@@
Distributionally robust optimization. Distributionally Robust Optimization (DRO) considers parameter uncertainty in optimization problems, and is particularly effective when only a limited number of uncertainty samples are available. The method compensates for the potential discrepancy between empirical and true uncertainty distributions by utilizing uncertainty descriptors like moment ambiguity sets (Van Parys et al., 2016), Kullback–Leibler ambiguity sets (Jiang and Guan, 2016), and Wasserstein ambiguity sets (Esfahani and Kuhn, 2018; Xie, 2021). The DRO framework has been increasingly utilized for its robust performance guarantees against distributional uncertainty in control (Boskos et al., 2024; Mestres et al., 2024; Long et al., 2023a; Chriat and Sun, 2024) and robotics (Ren and Majumdar, 2022; Coulson et al., 2021) applications.
@@
Lathrop et al. (2021) developed a Wasserstein safe variant of RRT, offering finite-sample probabilistic safety guarantees. A DRO approach was introduced by Ren and Majumdar (2022) to enhance policy robustness by iteratively training with adversarial environments generated through a learned generative model. Long et al. (2023b) proposed a DRO formulation for safe stabilizing control under model uncertainty, assuming that nominal safety and stability certificates are provided. Hakobyan and Yang (2022) introduced a distributionally robust risk map as a safety specification tool for mobile robots, reformulating optimal control problem over an infinite-dimensional probability distribution space into a tractable semidefinite program. Boskos et al. (2023) proposed a distributionally robust coverage control algorithm for a team of robots to optimally deploy in region with an unknown event probability density. A Wasserstein tube MPC is presented in Aolaritei et al. (2023) for stochastic systems, which utilizes Wasserstein ambiguity sets to construct tubes around nominal trajectories, enhancing robustness and efficiency with limited noise samples available. Compared to existing applications of DRO in robotics, our approach focuses on integrating DRO with control barrier functions to derive robust, sample-based safety constraints, enabling real-time navigation directly from sensor observations and estimated robot states in dynamic, unknown environments.
@@
Safe stabilizing control. Quadratic programming that integrates CLF and CBF constraints offers an efficient approach for synthesizing safe stabilizing control inputs for multi-robot navigation (Zhang et al., 2023), legged locomotion (Grandia et al., 2021), and humanoid operation (Khazoom et al., 2022). Despite the efficiency of the CLF-CBF QP in synthesizing safe stabilizing controls, it typically relies on perfect knowledge of system dynamics, state estimates, and barrier function constraints. However, in many robotics applications, various sources of uncertainty can significantly impact performance and reliability. Some recent studies have begun to explore this area, to address uncertainty in system dynamics (Dhiman et al., 2023; Emam et al., 2022), state estimates (Daş and Murray, 2022; Wang and Xu, 2023), and barrier function constraints (Long et al., 2021; Hamdipoor et al., 2023). These approaches typically utilize robust and probabilistic models to integrate uncertainty into the QP formulation, leading to convex reformulations that enhance robustness.
@@
Particularly in safe robot navigation, a robot might need to estimate barrier functions based on its (noisy) observations. Long et al. (2021) introduced an incremental online learning approach for estimating the barrier function from LiDAR data and proposed a robust reformulation of the CLF-CBF QP by incorporating estimation errors. Majd et al. (2021) integrate time-based rapidly-exploring random trees (RRTs) with CBFs to enable safe navigation in dynamic environments densely populated by pedestrians. Dawson et al. (2022) developed an approach to learn observation-space CBFs utilizing distance data and proposed a two-mode hybrid controller to avoid deadlocks in navigation. A reactive planning algorithm was presented in Liu et al. (2023b) for the safe operation of a bipedal robot with multiple obstacles, utilizing a single differentiable CBF derived from LiDAR point clouds. Abdi et al. (2023) proposed a method for learning vision-based CBF from RGB-D images with pre-training, enabling safe navigation of autonomous vehicles in unseen environments. By effectively decomposing and predicting the spatial interactions of multiple obstacles, Yu et al. (2023) proposed compositional learning of sequential CBFs, enabling obstacle avoidance in dense dynamic environments. Keyumarsi et al. (2024) introduced an efficient Gaussian Process-based method for synthesizing CBFs from LiDAR data, showcasing its effectiveness in a turtlebot navigation task. Zhang et al. (2024) introduced an efficient LiDAR-based framework for goal-seeking and exploration of mobile robots in dynamic environments, utilizing minimum bounding ellipses to represent obstacles and Kalman filters to estimate their velocities.
@@
Conformal Prediction. A related line of research to our work is the application of conformal prediction in robot navigation tasks. Conformal prediction (Shafer and Vovk, 2008; Zhao et al., 2024) is a statistical tool for uncertainty quantification that provides valid prediction regions with a user-specified risk tolerance, making it particularly useful for ensuring safety in dynamic environments. Several recent works have explored the integration of conformal prediction into motion planning and control frameworks. Lindemann et al. (2023) used conformal prediction to obtain prediction regions for a model predictive controller. Yang et al. (2023) employed conformal prediction to quantify state estimation uncertainty and design a robust CBF controller based on the estimated uncertainty. An adaptive conformal prediction algorithm was developed by Dixit et al. (2023) to dynamically quantify prediction uncertainty and plan probabilistically safe paths around dynamic agents.
3 Preliminaries
@@ 3.2 CLF-CBF Quadratic Program
where $\delta$ denotes a slack variable that relaxes the CLF constraint to ensure feasibility of the QP, controlled by the scaling factor $\lambda>0$. As discussed in Ames et al. (2017); Mestres et al. (2023), if the CBF $h$ has relative degree of $1$ with respect to the system dynamics in (1), the controller in (6) is Lipschitz continuous in $\mathbf{x}$ and piecewise continuous in $t$. This guarantees unique solutions for the closed-loop system, ensuring the safe set ${\cal C}(t)$ remains forward invariant. However, a few practical limitations should be noted: (a) the use of the slack variable $\delta$ relaxes the strict satisfaction of the CLF stability constraint, creating a trade-off between safety and stability that depends on the design of the safe set; (b) in numerical implementations, the CBF-QP must be solved at a fixed rate, whereas the theoretical guarantees are valid for executions in continuous time; and (c) while the CBF-QP is generally feasible, it may become infeasible if input constraints limit the available control actions. Researchers have proposed methods to address these challenges, including employing Input-Constrained Control Barrier Functions (ICCBFs) (Agrawal and Panagou, 2021), using neural network-based CBFs to account for input saturation (Liu et al., 2023c), and tuning of the class $\mathcal{K}_{\infty}$ functions (Parwana et al., 2022).